# Comparative Evaluation of 3D Reconstruction Methods for Object Pose Estimation

### v1

@@ A.2 Data Release
License.  The collected images and the reconstructed meshes are released under the CC BY 4.0 license: the license is conditioned on the license of the YCB [17] objects depicted in the images. The YCB objects are released under the Creative Commons Attribution 4.0 International (CC BY 4.0)  [^5] http://ycb-benchmarks.s3-website-us-east-1.amazonaws.com/, so the data is released under the same license.
@@ Appendix B Evaluation Setup
In the rest of this section, we report the implementation details relative to the evaluated 3D reconstructions (Sec B.1), the pose estimators used for the evaluation (Sec B.2), the pose evaluation metrics (Sec. B.3), and the nature of the objects in the evaluation dataset (Sec B.4).
B.1 Reconstruction Implementation Details
@@
B.2 Pose Estimators: FoundationPose and Megapose
@@ B.3 BOP Evaluation
$e_{\text{MSSD}}(\hat{\mathbf{P}},\bar{\mathbf{P}},S_{M},V_{M})=\text{min}_{\mathbf{S}\in S_{M}}\text{max}_{\mathbf{x}\in V_{M}}\|\hat{\mathbf{P}}\mathbf{x}-\bar{\mathbf{P}}\mathbf{S}\mathbf{x}\|_{2}\enspace.$ (1)
@@
$e_{\text{MSPD}}(\hat{\mathbf{P}},\bar{\mathbf{P}},S_{M},V_{M})=\\ \text{min}_{\mathbf{S}\in S_{M}}\text{max}_{\mathbf{x}\in V_{M}}\|\text{proj}(\hat{\mathbf{P}}\mathbf{x})-\text{proj}(\bar{\mathbf{P}}\mathbf{S}\mathbf{x})\|_{2}\enspace.$ (2)
@@
$e_{\text{VSD}}(\hat{D},\bar{D},\hat{V},\bar{V},\tau)=\mathrm{avg}_{p\epsilon\hat{V}\cup\bar{V}}\\ \begin{cases}0&\text{if}~{}p\in\hat{V}\cap\bar{V}\wedge\|\hat{D}(p)-\bar{D}(p)\|\leq\tau\\ 1&\text{otherwise }\end{cases}\enspace.$ (3)
@@
B.4 Object Categorization
@@
Appendix C Additional Results
We complete the experimental results of the main paper with the average recall on each of the three pose errors (Sec. C.1), the influence of the size of the training data on the performance (Sec. C.2), further results on the relation between the object categories and the pose performance (Sec. C.3), the importance of the texture for pose estimation (Sec C.4), the detail of the pose performances per objects (Sec C.5), and qualitative results showing renderings from the reconstructed meshes (Sec C.6).
C.1 Pose Evaluation: Average Recall
@@
The BOP benchmark defines multiple errors relevant to assess how suitable a reconstruction is for a given application such as augmented and virtual reality (MSPD), navigation (VSD), and object manipulation (MSSD). In the main paper, the pose performances are measured with the Average Recall (AR) averaged over all three VSD, MSSD, and MSPD errors We now report the AR for each error separately (Fig. 7) to support the conclusions of the main paper (Sec.5 "Finer Pose Evaluation").
@@
The main paper mentions the performance drop induced when replacing the original CAD models with the reconstructed ones and how it is consitent between the two pose estimators for the best reconstruction methods, i.e., RealityCapture [82], Nerfacto [100], VolSDF [113], MonoSDF [116], BakedSDF [114], and NeuralAngelo [59]. This remains the case for the individual ARs and we also observe that the relative performances of the reconstruction methods are consistent across the different errors. So one can expect that improving the 3D reconstructions would benefit all three pose estimation errors, hence various applications involving object pose estimation.
@@
C.2 Reconstruction with varying training size
@@
C.3 Object Categories
We split the YCB-V [111] objects into 8 groups defined in Sec. B.4 based on their shape, texture, material properties, and the degree of change between the collected objects and the original ones. We report results for all object categories in Fig. 9 (including the 3 reported in the main paper) and observe that the relative pose performance across object categories is telling of some of the current challenges in 3D reconstruction.
@@ C.3 Object Categories
C.4 Influence of the Texturing Algorithm
@@
C.5 Per-Object Pose Evaluation

### v2

@@ 7.2 Data Release
License.  The collected images and the reconstructed meshes are released under the CC BY 4.0 license: the license is conditioned on the license of the YCB [15] objects depicted in the images. The YCB objects are released under the Creative Commons Attribution 4.0 International (CC BY 4.0)   [^6] http://ycb-benchmarks.s3-website-us-east-1.amazonaws.com/ so the data is released under the same license.
@@ 8 Evaluation Setup
In the rest of this section, we report the implementation details relative to the evaluated 3D reconstructions (Sec 8.1), the pose estimators used for the evaluation (Sec 8.2), the pose evaluation metrics (Sec. 8.3), and the nature of the objects in the evaluation dataset (Sec 8.4).
8.1 Reconstruction Implementation Details
@@
8.2 Pose Estimators: FoundationPose and Megapose
@@ 8.3 BOP Evaluation
$e_{\text{MSSD}}(\hat{\mathbf{P}},\bar{\mathbf{P}},S_{M},V_{M})=\\ \text{min}_{\mathbf{S}\in S_{M}}\text{max}_{\mathbf{x}\in V_{M}}\|\hat{\mathbf{P}}\mathbf{x}-\bar{\mathbf{P}}\mathbf{S}\mathbf{x}\|_{2}\enspace.$
@@
$e_{\text{MSPD}}(\hat{\mathbf{P}},\bar{\mathbf{P}},S_{M},V_{M})=\\ \text{min}_{\mathbf{S}\in S_{M}}\text{max}_{\mathbf{x}\in V_{M}}\|\text{proj}(\hat{\mathbf{P}}\mathbf{x})-\text{proj}(\bar{\mathbf{P}}\mathbf{S}\mathbf{x})\|_{2}\enspace.$
@@
$e_{\text{VSD}}(\hat{D},\bar{D},\hat{V},\bar{V},\tau)=\mathrm{avg}_{p\epsilon\hat{V}\cup\bar{V}}\\ \begin{cases}0&\text{if}~{}p\in\hat{V}\cap\bar{V}\wedge\|\hat{D}(p)-\bar{D}(p)\|\leq\tau\\ 1&\text{otherwise }\end{cases}\enspace.$
8.4 Object Categorization
@@
9 Additional Results
We report the quantitative results that support the conclusions drawn in the main paper and that could not fit due to the page limit: the average recall on each of the three pose errors (Sec. 9.1), the influence of the size of the training data on the performance (Sec. 9.2), further results on the relation between the object categories and the pose performance (Sec. 9.3), the importance of the texture for pose estimation (Sec 9.4), the detail of the pose performances per objects (Sec 9.5), and qualitative results showing renderings from the reconstructed meshes (Sec 9.6).
9.1 Pose Evaluation: Average Recall
The BOP benchmark defines multiple errors relevant to assess how suitable a reconstruction is for a given application such as augmented and virtual reality (MSPD), navigation (VSD), and object manipulation (MSSD) [45, 46, 2]. In the main paper, the pose performances are measured with the Average Recall (AR) averaged over all three VSD, MSSD, and MSPD errors We now report the AR for each error separately (Fig. 7) to support the conclusions of the main paper (Sec.5 "Finer Pose Evaluation").
@@ 9.1 Pose Evaluation: Average Recall
The main paper mentions the performance drop induced when replacing the original CAD models with the reconstructed ones and how it is consitent between the two pose estimators for the best reconstruction methods, i.e., RealityCapture [82], Nerfacto [100], VolSDF [113], MonoSDF [117], BakedSDF [114], and NeuralAngelo [58]. This remains the case for the individual ARs and we also observe that the relative performances of the reconstruction methods are consistent across the different errors. So one can expect that improving the 3D reconstructions might benefit all three pose estimation errors, hence various applications involving object pose estimation.
@@
9.2 Reconstruction with varying training size
@@
9.3 Object Categories
We split the YCB-V [112] objects into 8 groups defined in Sec. 8.4 based on their shape, texture, material properties, and the degree of change between the collected objects and the original ones. We report results for all object categories in Fig. 9 (including the 3 reported in the main paper) and observe that the relative pose performance across object categories is telling of some of the current challenges in 3D reconstruction.
@@ 9.3 Object Categories
9.4 Influence of the Texturing Algorithm
@@
9.5 Per-Object Pose Evaluation